Milden Company is a distributor who wants to start using a contribution format income statement for planning purposes. The company has analyzed its expenses and developed the following cost formulas:

Cost Cost Formula
Cost of good sold $27 per unit sold
Advertising expense $177,000 per quarter
Sales commissions 6% of sales
Shipping expense ?
Administrative salaries $87,000 per quarter
Insurance expense $9,700 per quarter
Depreciation expense $57,000 per quarter

Because shipping expense is a mixed cost, the company needs to estimate the variable shipping expense per unit sold and the fixed shipping expense per quarter using the following data:

Quarter Units Sold Shipping
Expense
Year 1:
First 23,000 $ 167,000
Second 25,000 $ 182,000
Third 30,000 $ 224,000
Fourth 26,000 $ 187,000
Year 2:
First 24,000 $ 177,000
Second 27,000 $ 192,000
Third 42,250 $ 244,000
Fourth 39,250 $ 220,000

Required:

1. Using the high-low method, estimate a cost formula for shipping expense in the form Y = a + bX.

2. In the first quarter of Year 3, the company plans to sell 29,000 units at a selling price of $57 per unit. Prepare a contribution format income statement for the quarter.
Answer #1
Variable cost p.u = Change in cost/change in activity
Variable cost p.u = 244000-167000/42250-23000
variable cost p.u = 4
fixed cost = total cost-total variable cost
fixed cost = 244000- (42250*4)
Fixed cost = 75000
cost formula for shipping exp = Y = 75000+4X
2) Contribution income statement
Particulars amount
sale (29000*57) 1653000
less: variable cost
cost of goods sold (29000*27) 783000
sale commission (1653000*6%) 99180
shipping expenses (29000*4) 116000
total variable cost 998180
contribution margin 654820
less: fixed cost
advertising exp 177000
administrative exp 87000
shipping expenses 75000
insurance exp 9700
depreciation expenses 57000
total fixed cost 405700
net operating income 249120

Give a thumbs up if it is helpful

please let me know if any doubt